OpenStudy (anonymous):

The measure of each interior angle of a regular polygon is 150°. How many sides does the polygon have? 15 10 12 8

OpenStudy (anonymous):

Where n is the number of sides, angle = 180 * (n-2) / n 150 * n = 180 * (n - 2) 150n = 180n - 360 150n + 360 = 180n 360 = 30n n = 12 This is a 12-sided polygon

OpenStudy (anonymous):

Okay thanks I get it.